[figure: 9.7 cm]

TUNNEL CONDUCTIVITY OF CUPRATES WITH NEGATIVE

MAGNETORESISTANCE

V.Yu. Tarenkov, V.M. Svistunov, Yu.V. Medvedev and S.J. Lewandowski^{a} Donetsk Physical-Technical Institute of the National Academy of Sciences of Ukraine,

72 R.Luxemburg, Donetsk, 340114, Ukraine aInstytut Fizyki, Polish Academy of Sciences, PL 02-668 Warszawa, Poland

We have performed tunneling experiments on ceramic LaSrMnO compounds exhibiting negative magnetoresistance. Two types of tunnel junctions were investigated: symmetric ceramic-insulator-ceramic and asymmetric ceramic-insulator-metal (lead). The conductance G of asymmetric junctions was very sensitive to the applied magnetic field H, both at null polarization V for temperatures T > T_{c}, where T_{c} denotes the superconducting transition temperature of lead, and at polarization beyond the gap voltage for T < T_{c}. Symmetric junctions of the ”break” type exhibited hysteretic G(H) dependence and 15% change of conductivity in applied magnetic field H=100 Oe. A characteristic feature of the observed characteristics is the presence of linear G(V) background, indicating that tunneling takes place through a system of localized states inside the barrier. The existence of such states is characteristic for tunnel barriers constituted by metal oxides.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Vladimir Tarenkov

Address for correspondence : Donetsk Physical-Technical Institute of NASU 72, str. R.Luxemburg 340114 Donetsk Ukraine FAX: (0622) 55-74-62

Email address : medvedev@kinetic.ac.donetsk.ua